OpenStudy (anonymous):

Solve the following system of equations. x + 3y + 2z = –16 2x – y + 2z = –15 2x – 2y – z = 1

OpenStudy (anonymous):

x=-3 y=-1 z=-5

OpenStudy (anonymous):

First step is eliminate the Z x + 3y + 2z = -16 2x - y + 2z = -15 ------------------------------ (-) -x + 4y = -1 ----------------------------------------… (1) x + 3y + 2z = -16 x + 3y + 2z = -16 2x - 2y - z = 1 x2 4x - 4y - 2z = 2 ---------------------------------- (+) 5x- y = - 14 ----------------------------------------… (2)

OpenStudy (anonymous):

Second step is eliminate the Y (1) and (2) -x + 4y = -1 -x + 4y = -1 5x - y = -14 x 4 20x - 4y = -56 ------------------------ (+) -19x = -57 x = 57/19 ----------------------------------------… (3) After get the X value, put in equation (2) to get the Y value 5x - y = -14 - y = - 14 -5x y = 14 + 5 x Next after you get the X and Y, put in your first equation x + 3y + 2z = -16 2z = -16 -x-3y z = -8 - 1/2 x - 3/2y then you get the Z.... sorry not much time to calculate the rest for you :)

OpenStudy (anonymous):

thanks you guys c: